OpenStudy (anonymous):

Solve for n: (4^n)^3 (2^n + 2) = (4^2)^n – 1

OpenStudy (anonymous):

Please anyone help!!!! :(

OpenStudy (anonymous):

try this, try solving the right side. do what's in the parenthesis first

OpenStudy (anonymous):

I tried but no success :(

OpenStudy (anonymous):

wait I think I figured something out I'll get back to you.

OpenStudy (anonymous):

wait is it 2^n plus 2 or 2^n divided by 2

OpenStudy (anonymous):

2^n plus 2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

could you write it out in equation please.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve for n: (4^n)^3 (2^n + 2) = (4^2)^n – 1.

OpenStudy (anonymous):

ok so\[(4^{n})^{3}(2^{n+2})=(4^{2})^{n+1}\]does that look right?

OpenStudy (anonymous):

if so I understand it and can walk you through it.

OpenStudy (anonymous):

Yes this is how it looks like! :)

OpenStudy (anonymous):

ok first thing, you'll have to trust me, this is dificult. change the 4s to \[2^2\]

OpenStudy (anonymous):

2^2s

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

yes. your equation should now look like this\[(2^{6n})(2^{n+2})=(2^{4})^{n-1}\]

OpenStudy (anonymous):

that's (2^(6n))(2^(n+2))=(2^4)^(n-1)

OpenStudy (anonymous):

I solved it!!!

OpenStudy (anonymous):

n=-2

OpenStudy (anonymous):

wow!! THANK you so much

OpenStudy (anonymous):

no problem, i had to ask my dad for help though so thank him.

OpenStudy (anonymous):

Thank you so so so so much!!!! And thank your dad for me so much!!! :)

OpenStudy (anonymous):

any time, what math are you in that was hard!

OpenStudy (anonymous):

Algebra 2 with trigonometry

OpenStudy (anonymous):

It is quite complicated sometimes :(

OpenStudy (anonymous):

I'm in algebra 2 also, but without the trig,( except for the last chapter we did, that was weird.)

OpenStudy (anonymous):

Ah cool, do you like math?

OpenStudy (anonymous):

I like it but sometimes I get so stuck :(

OpenStudy (anonymous):

I admit, i'm a nerd. i LOVE it, but it runs in my blood you're really good, especially at problem solving.

OpenStudy (anonymous):

Wow, you're lucky that it runs in your blood!!! Thank you but I think I am not that good at it :( but thanks!!! for everything!

OpenStudy (anonymous):

any time